### v1

@@ 7.5.2 Subpopulation shift
Subpopulation shift refers to changes in the proportion of some subpopulations between training and deployment [221]. In these contexts, DL models may have high overall performance yet still underperform in rare subgroups. Subpopulations shift can be categorized into spurious correlations, attribute imbalance, class imbalance and attribute generalization. Spurious correlations refers to non-causal relationships between the input and the label that may shift in deployment (non-robust features, namely, image backgrounds or texture). Another type of subpopulation shift is attribute imbalance in which certain attributes are sampled with a much smaller probability than others in the training. Class imbalance occurs when class labels exhibit imbalanced distributions, which may cause lower preference for minority labels. Attribute generalization denotes the setting where some attribute are totally missing in the training domain but present in the testing domain.
@@ 7.5.4 Continual DG
Conventional DG assumes that multiple source domains are accessible and the domain shift is abrupt. However, this is not universally applicable to all real-world applications where the data distribution may gradually change over time, especially, in the medical field. In this context, new disease or new biomarkers may arise. As the domain continues to evolve, new domains will consistently emerge. Re-training DL models, under the conventional scheme of DG, to keep-to-date with both new and existing domains can be both resource-intensive and inefficient. Continual learning, on the other hand, permits the model to continuously learn from a sequence of tasks over time while maintaining performances on all experienced tasks. Combining continual learning and DG would enable to model the evolutionary patterns of temporal domains and leveraging these patterns to palliate the distribution shift in the future domains. Recent work [211] proposed a continual domain generalization over temporal drifts, where the goal is to generalize on new unseen domain given that only data from the current domain is accessible at any given time, while information from past domains is unavailable.
@@ 7.5.5 Unified Benchmarking
In the benchmarking framework, researcher could also state what problem of shift they are targeting to solve. In this work, we have identified two main categories of methods based on the shift type: covariate and concept shift. However, in the medical field, the domain shift problem is very complex and both shifts can appear simultaneously. While it is more challenging to tackle both problems, future work handling full shift (covariate shift and concept shift) holds great potential for the clinical world.
@@ Conclusion
In the medical field, data exhibit different sources of variation: images may be collected from multiple countries and different ethnic group (causing covariate shift), data can be gathered using different criteria (different screening programs), annotations differences, etc. (causing concept shift). To mitigate these challenges, we reviewed state-of-the-art methods for the generalization of DL models in medical image classification and discussed challenges and future research trends for this line of research. We hope that this work will help the research community to tackle the problem of generalization in a variety of applications.

### v3

@@ 7.4.1 Datasets in DG
7.4.2 Modelling pipeline strategies
The majority of the work in generalization research tackled the train-test data shift (also called train-test locus), i.e., considering the classical shift type between train and test data. Other types of shift loci can be investigated as proposed by Hupkes et al. [14]. Namely, the fine-tune train test locus which refers to the situation where a model is evaluated on a finetuning test set that has a different distribution from the finetuning training data. In this context, finetuning could be achieved by refining all the model parameters, freezing the network’s top layer and training only the dense layers [220], or a few of the final convolutional layers [221]. Another type is the pretrain-train locus which evaluates if a specific pretraining method produces models that are effective when subsequently trained on diverse tasks or domains. This is often evaluated in the case of foundation models. The pre-train-test locus is encountered when a pre-trained model is tested directly on out-of-domain data.
@@ 7.4.3 Subpopulation shift
Biases in DL models, associated with factors such as race, gender or age can result in healthcare disparities and negative patient outcomes. In fact, underrepresented training data can lead to subpotimal DL models. One key contributing factor to this is subpopulation shift, i.e, changes in the proportion of some subpopulations between training and deployment [222]. In these contexts, DL models may have high overall performance yet still underperform in rare subgroups. Subpopulations shift can be categorized into spurious correlations, attribute imbalance, class imbalance and attribute generalization. Spurious correlations involve non-causal relationships between the input and the label that may shift during deployment, such as image backgrounds or texture. Attribute imbalance occurs when certain attributes are sampled with a much smaller probability than others in the training. Class imbalance happens when class labels are distributed unevenly, leading to lower preference for minority labels. Attribute generalization refers to the setting where some attributes are absent in the training domain but present in the testing domain.
@@ 7.4.5 Continual DG
Conventional DG assumes that multiple source domains are accessible and the domain shift is abrupt. However, this is not universally applicable to all real-world applications where the data distribution may gradually change over time, especially, in the medical field. In this context, new disease or new biomarkers may arise. As the domain continues to evolve, new domains will consistently emerge. Re-training DL models, under the conventional scheme of DG, to keep-to-date with both new and existing domains can be both resource-intensive and inefficient. While the transfer learning paradigm seems to be an effective strategy to solve this problem, it should be carefully applied to DG models. For instance, Garrucho et al. [32] demonstrated that fine-tuning a DG model to unseen domain can sometimes decrease performance. In the medical domain, transfer learning faces challenges such as data availability and catastrophic forgetting. Fine-tuning models in new domains can lead to overfitting to less diverse datasets and forgetting previously learned information. This could be attributed to a small dataset or even noisy data. Samala et al. [225] noted that training with noisy data, even with as few as 10% corrupted labels, could increase generalization error. Therefore, it is also not recommended to perform transfer learning when the quality of the data is poor.
A potential future direction to address these challenges is continual learning. It permits the model to continuously learn from a sequence of tasks over time while maintaining performances on all experienced tasks. Combining continual learning and DG would enable to model the evolutionary patterns of temporal domains and leveraging these patterns to palliate the distribution shift in the future domains. Recent work [226] proposed a continual domain generalization over temporal drifts, where the goal is to generalize on new unseen domain given that only data from the current domain is accessible at any given time, while information from past domains is unavailable.
@@ 7.4.6 Unified Benchmarking
In the last few years, benchmarking has shown a great interest in the medical research community. In general, DG performances are compared to a baseline approach, Empirical Risk Minimization (ERM), where a single model is learned on pooled data across all training sources by minimizing the global average risk. Several applications were targeted, Zhang et al. [229] benchmarked [^40] https://github.com/MLforHealth/ClinicalDG the performance of eight DG methods on multi-site clinical times series from Intensive Care Units (ICUs) and chest X-ray imaging data from four sites. In line with prior work on general imaging datasets [213], their experiments on real-world medical imaging data revealed that the current DG methods do not consistently achieve significant gains in OOD performance over ERM. More recently, Che et al. [230] targeted DR grading in unseen domains. They presented a unified framework named Generalizable Diabetic Retinopathy Grading Network  [^41] https://github.com/chehx/DGDR, which demonstrated promising performances compared to ERM. In addition, for fair evaluations, they have provided a publicly available benchmark, the GDR-Bench Dataset, which includes eight open-source fundus datasets. In line with this study, future work should aim to propose real-world benchmark datasets for different medical modalities specifically for DG. This initiative would undoubtedly promote standardized evaluation protocols, ensuring consistency and reliability in the assessment of DG methods.
7.4.7 Privacy concerns
Learning under domain shift in the medical field is also subject to data privacy and regulatory concerns. In certain cases, it is challenging for a single institution to collect enough diverse data, especially for rare diseases. Multi-DG in these settings can facilitate data collection from multiple institutions, aiming to develop models that generalize to unseen domains. Integrating federated learning to DG is a promising solution to ensure data privacy and compliance with regulations, enabling collaborative efforts without sharing sensitive patient data directly. This approach not only preserves patient confidentiality but also ensures that collaborative research adheres to legal standards such as data protection laws GDPR and HIPPA.
7.4.8 Multimodal DG
A promising research direction involves integrating medical multimodality into DG. Multimodality encompasses combining various data types such as electronic health records, imaging techniques including 2D and 3D image information [231], and genomic data. This integration adds complexity to assessing the dataset shift. For instance, there might be scenarios where data from one modality is missing, some data is noisy, unannotated, has unreliable labels, or is scarce during the training or testing phases. Recent work has targeted to solve these problems. For example, generating missing data using generative models. This, however, may exacerbate the problem by possibly introducing a generated shift. Despite these challenges, this area of research is crucial as it emulates the comprehensive diagnostic methodology employed by medical professionals, and allows for improved DL performances. Indeed, a future direction involves developing innovative multimodal and multidomain AI models for clinical decision-making using foundation models.
7.5 Limitations
One limitation of this work is that we only considered Scopus as a database, which may not be representative of all existing work done in this field. Another limitation is that while we focused on two main shifts, we acknowledge that domain shift is more complex in the medical field. These assumed shift (covariate/concept shifts) assume that one of the probability distribution is fixed. However, in real scenarios, this may also be more complex and both shifts can appear simultaneously. While it is more challenging to tackle both problems, future work handling full shift (covariate shift and concept shift) holds great potential for the clinical world. We note that there is a limited consensus on the terminology used in papers. Shift types are defined differently in some papers and new terms can arise as acquisition shift [32]. This make the search suboptimal for literature review. A unified terminology as proposed in our work and [16] would help researchers to rely on a unifying framework for addressing domain shift.
@@ Conclusion
In the medical field, data exhibit different sources of variation: images may be collected from multiple countries and different ethnic group (causing covariate shift), data can be gathered using different criteria (different screening programs), annotations differences, etc. (causing concept shift). To mitigate these challenges, we reviewed state-of-the-art methods for the generalization of DL models in medical image classification and discussed challenges and future research trends for this line of research. We hope that this work will help the research community to tackle the problem of generalization in a variety of applications. Beyond out-of-domain generalization, achieving a fully trustworthy and responsible model in healthcare requires robustness against malicious (adversarial) attacks, and interpretability. Securing both the data and the models is crucial, especially in medical diagnosis and clinical settings, given the growing regulatory concerns. Interpretability allows for understanding how a model makes its predictions and assessing their validity, which builds trust in the model and ensures appropriate use. In addition, for safe deployment in real world clinical applications, AI models must express uncertainty when operating outside their training data range. Ultimately, the methods discussed in this survey could democratize access to AI by offering a scalable screening, more reliable diagnosis and more equitable access to high quality care. Robust clinical validation across various institutions and demographics would further promote the wider adoption of AI in healthcare.
Acknowledgement
This work was supported by the French National Research Agency under the LabCom program (ANR-19-LCV2-0005 - ADMIRE project).
Declaration of competing interests
The authors declare that there are no known competing interests that could affect this work.
Registration and protocol
The review was not registered. A review protocol was not prepared for this systematic review.